### v1

@@ Abstract
Numerous Deep Learning (DL) models have been developed for a large spectrum of medical image analysis applications, which promises to reshape various facets of medical practice. Despite early advances in DL model validation and implementation, which encourage healthcare institutions to adopt them, some fundamental questions remain: are the DL models capable of generalizing? What causes a drop in DL model performances? How to overcome the DL model performance drop? Medical data are dynamic and prone to domain shift, due to multiple factors such as updates to medical equipment, new imaging workflow, and shifts in patient demographics or populations can induce this drift over time. In this paper, we review recent developments in generalization methods for DL-based classification models. We also discuss future challenges, including the need for improved evaluation protocols and benchmarks, and envisioned future developments to achieve robust, generalized models for medical image classification.
@@
[…]
@@ Introduction
Deep Learning (DL) models are the current state-of-art method for medical image classification. The availability of high quality labeled data, typically through multi-site collaboration projects, has paved the way to employ these data driven-based approaches in supervised medical image analysis. Nowadays, DL models have achieved human level performances in different medical domains such as dermatology [55], oncology [127], histopathology [77] and ophthalmology [45].
@@
However, UDA methods are limited in practice, as they still require access to a part of the test-domain data during the training procedure. To overcome this limitation, Domain Generalization (DG) methods have emerged as a more promising solution. In DG, the goal is to develop a DL model that is able to generalize to one unseen target domain via learning from a single or multiple source domains, without having access to the testing data from the target domain. However, training DG methods using multi-source data (multi-DG) has been considered as costly since collecting medical from multiple sources is challenging, and medical data are subject to privacy regulations. To address this problem, recent work focused on an additional research line, called single domain generalization (single-DG), in which the goal is to develop a DL model that is able to generalize to multiple target domains via learning from a single source domain [115]. Alternatively, semi-supervised domain generalization [231] combines the single-DG and multi-DG by using one labeled sources domain and several unlabeled source domain to boost the performances.
@@
DG in computer vision dataset is becoming an emerging field: numerous surveys have been proposed [235, 204]. In the medical field, research has focused on domain adaptation [67] or unsupervised domain adaptation [101]. Other medical research has reviewed the problem of learning with noisy labels [93, 238, 153]. However, to the best of our knowledge, no medical review has studied the problem of generalization of DL models in the medical field with a focus on both domain shift problems: covariate shift and concept shift. This paper presents the first systematic review of generalization research in medical image classification.
For the literature review, we used the following query on Scopus: “domain generalization”  OR “noisy labels” OR “covariate shift” OR “concept shift”. We included papers published from 1 January 2020 to 10 April 2023 (included). A total of 2086 papers were found. First, the search results were reviewed and duplicated records were removed. This resulted in 2027 papers. Abstracts and titles were manually reviewed and papers which are in the medical field were retained. A total of 372 papers were selected. Then, we selected the papers dealing with medical image classification and deep learning methods. Thus, 144 papers were selected. Out of these papers, 3 papers were excluded because they were not in English, and 3 other papers were excluded because they were reviews (these reviews [238, 93, 153] were previously introduced in this section). Furthermore, one paper was excluded as it presented the results of a challenge. Therefore, 137 articles were retained. For the purpose of assessing DL generalizability, using more than one dataset is desirable. We considered papers including more than one dataset, or including a multi-center dataset. While the main focus was on DG, we did not restrict our query to DG methods (or a type of DG, i.e., multi-DG) as UDA methods can also be extended to DG. This corresponds to 82 articles. Among these papers, 4 papers [125, 162, 197, 99] did not propose a new method for tackling concept shift or covariate shift, so they were discarded. Thus, the total number of papers considered in this study was 78 papers.
@@
The organization of this paper is as follows. In Section 3, we briefly describe the problem of domain generalization. In Section 4, we present public medical datasets used for generalization research. Section 5 summarizes previous findings in clinical domain generalization benchmarks. In Section 6, we present our taxonomy, in which we review DL methods that have dealt with covariate shift in the medical domain (Section 6.1) and DL methods aiming to overcome the problem of concept shift and noisy labels (Section 6.2). Section 7 discusses the benefit of current DG methods based on the results of challenge data. Furthermore, it presents trends in DG development, related research to DG, implementation details, and future directions. Finally, Section 8 concludes this work.
@@
Public medical datasets for generalization research
There exist many public datasets which have been adopted for generalization research in the medical field. Some of these datasets were proposed as part of a challenge. Table 1 and Table 2 summarize the public medical datasets which can be used for generalization research. In this section, we will give brief details for public datasets available as part of a challenge and which are the most relevant in DG context.
4.0.0.1 MIDOG datasets
Mitosis Domain Generalization (MIDOG) dataset targets the detection of mitotic figures in histopathology images under domain shift regime.
  • MIDOG 2021 dataset [12]: This dataset was part of the MICCAI MIDOG 2021 challenge which aims to evaluate methods that mitigate domain shift and derive scanner-agnostic algorithms. It addresses DG in histopathology. The main task was mitosis detection in breast cancer. The challenge dataset features 300 cases, 6 scanners, and more than 2500 mitosis. The domains are defined by scanner types. Namely, Hamamatsu NanoZoomer XR (C12000-22, Hamamatsu, Hamamatsu City, Japan) (Scanner A), Hamamatsu NanoZoomer S360 (Hamamatsu, Hamamatsu City, Japan) (Scanner B), Aperio Scanscope CS2 (Leica Biosystems, Nussloch, Germany) (Scanner C) and Leica Aperio GT 450 (Leica Biosystems, Nussloch, Germany)(Scanner D), 3DHISTECH Panoramic 1000 (3DHISTECH, Budapest, Hungary) (Scanner E), Hamamatsu NanoZoomer 2.0RS (C10730-12, Hamamatsu, Hamamatsu City, Japan) (Scanner F).
  • MIDOG 2022 dataset: It originates from MIDOG 2022 challenge [11]. It is a multi-tumor, multi-species, multi-laboratory and multi-scanner dataset. The main task was mitosis detection. The goal of this challenge is to develop strategies that lead to machine learning solutions that are invariant to this tissue-related domain-shift. 6 tumor domains were provided for developing the DL models: Human breast carcinoma, retrieved from the University Medical Center Utrecht (Domain A), Canine lung carcinoma, retrieved from the University of Veterinary Medicine Vienna (VMU) (Domain B), Canine lymphosarcoma, retrieved from the VMU (domain C), Canine cutaneous mast cell tumor, retrieved from the DA of FUB (Domain D), Human pancreatic and gastrointestinal neuroendocrine tumor, retrieved from the UMC Utrecht (Domain E). The evaluation was carried out on ten independent tumor domains.
  • MIDOG++ dataset [13] is an extension of MIDOG 2021 and 2022 challenge datasets. It consists of 7 tumor types: breast carcinoma, lung carcinoma, lymphosarcoma, cutaneous mast cell tumor, neuroendocrine tumor, soft tissue sarcoma, and melanoma. The domains of the dataset include different tumor types, laboratories, whole slide image scanners, and species.
4.0.0.2 CAMELYON datasets
Cancer Metastases in Lymph nodes challenge (CAMELYON) datasets target the automated detection of cancer metastases in Whole-Slide Images (WSIs) of sentinel lymph nodes.
4.0.0.3 UBC-OCEAN dataset
The University of British Columbia - Ovarian Cancer Subtype Classification and Outlier Detection (UBC-OCEAN) competition aims to classify ovarian cancer subtypes from microscopy scans of biopsy samples  [^1] https://www.kaggle.com/competitions/UBC-OCEAN/overview. The UBC-OCEAN dataset consists of an extensive ovarian cancer dataset of histopathology images obtained from more than 20 medical centers. However, the origin of the image is not given in the meta-data. The test set contains images from different source hospitals than the train set. The class other is not present in the training set.
4.0.0.4 LUNA-16
: The goal of LUNA-16 challenge is the automated detection of pulmonary nodules in thoracic Computed Tomography (CT) scans [169]. This challenge use data from a large public LIDC-IDRI dataset  [9]. More precisely, scans with a slice thickness greater than 2.5 mm were excluded. The resulting dataset contains 888 CT scans.
4.0.0.5 PANDA dataset
: The goal of Prostate Cancer Grade Assessment (PANDA) challenge [27] is the diagnosis of prostate cancer in biopsies. It aims to develop AI algorithms for Gleanson grading. In total, the PANDA dataset comprises 12,625 WSIs of prostate biopsies retrospectively collected from 6 different sites for algorithm development, tuning and independent validation. Cases for development, tuning and internal validation originated from two European (EU) centers: Radboud University Medical Center, Nijmegen, the Netherlands and Karolinska Institutet, Stockholm, Sweden. The external validation data consisted of a US (741 cases) and an EU set (330 cases).
Dataset Modality (Organs) Number of cases Reference MIDOG 2021 dataset [12] Histopathology (Breast) 300 images https://midog2021.grand-challenge.org/ VGH [19] 5,920 images https://tma.im/tma_portal/C-Path/supp.html NKI [19] 8,337 images Camelyon16 WILDS dataset [20] 399 WSI https://camelyon16.grand-challenge.org/Data/ PatchCamelyon [196] 327,680 color images https://patchcamelyon.grand-challenge.org/ Camelyon17 WILDS dataset [100] 1000 WSI https://camelyon17.grand-challenge.org/Data/ TCGA-BRCA 1,098 cases https://portal.gdc.cancer.gov/projects/TCGA-BRCA BACH dataset Microscopy,Histopathology (Breast) 400 microscopy images 30 WSI https://iciar2018-challenge.grand-challenge.org/Dataset/ MIDOG 2022 dataset [11] Histopathology (Multi-organs) 405 images https://midog2022.grand-challenge.org/ MIDOG++ dataset [13] Histopathology (Multi-organs) 503 images https://github.com/DeepMicroscopy/MIDOGpp UBC-OCEAN Histopathology (Ovaries) 538 training images https://www.kaggle.com/competitions/UBC-OCEAN/data PANDA Histopathology (Prostate) 12,625 WSIs https://www.kaggle.com/c/prostate-cancer-grade-assessment DiagSet-B 4675 scans https://github.com/michalkoziarski/DiagSet SICAPv2 [179] 155 biopsies (95 patients) https://data.mendeley.com/datasets/9xxm58dvs3/1 TUPAC-16[22] Histopathology (Colon) 1076 cases https://tupac.grand-challenge.org/Dataset/ Kather16 [95] Histopathology (Colon) 5,000 patches https://zenodo.org/records/53169 Kather19 [94] 100,000 patches http://dx.doi.org/10.5281/zenodo.1214456 CRC-TP [89] 196,000 patches https://warwick.ac.uk/TIAlab/data/crchistolabelednucleihe/. IHC [120] 1,376 images http://fimm.webmicroscope.net/supplements/epistroma CRC-VALHE-7K 7,180 image patches https://zenodo.org/records/1214456 Stanford-CRC [220] 66,578 image tiles https://github.com/rikiyay/MSINet CRC-DX-TRAIN dataset 93,408 image tiles https://zenodo.org/records/2530835#.XwCkDZNKhTY CRC-DX-TEST dataset 99,904 image tiles https://zenodo.org/records/2530835#.XwCkDZNKhTY Chaoyang Dataset [241] 6,160 WSI https://bupt-ai-cz.github.io/HSA-NRL/ DigestPath2019 [113] 690 patients https://digestpath2019.grand-challenge.org/ TCGA-LGG Histopathology (Brain) 516 cases https://portal.gdc.cancer.gov/projects/TCGA-LGG TCGA-GBM 617 cases https://portal.gdc.cancer.gov/projects/TCGA-LGG TCGA-LUAD Histopathology (Lung) 585 cases https://portal.gdc.cancer.gov/projects/TCGA-LUAD TCGA-LUSC 504 cases https://portal.gdc.cancer.gov/projects/TCGA-LUSC ADNI-1 [87] MRIs (Brain) 748 subjects https://adni.loni.usc.edu/data-samples/access-data/ ADNI-2 [87] 708 subjects https://adni.loni.usc.edu/data-samples/access-data/ AIBL [54] 549 subjects https://adni.loni.usc.edu/data-samples/access-data/ ISIC-2017 [40] Dermoscopic images 2,750 images https://challenge.isic-archive.com/data/ ISIC-2019[193] 33,569 images HAM10000 (HAM)[193] 10,000 images https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/DBW86T Dermofit (DMF) 1,300images https://homepages.inf.ed.ac.uk/rbf/DERMOFIT/datasets.htm Derm7pt (D7P) [96] 2,000 images http://derm.cs.sfu.ca INbreast [135] Mammography (Breast) 115 cases https://www.kaggle.com/datasets/ramanathansp20/inbreast-dataset OPTIMAM dataset [73] 179,326 cases https://medphys.royalsurrey.nhs.uk/omidb/ BCDR [124] 3,703 digitised film mammograms https://www.medicmind.tech/cancer-imaging-data LIDC-IDRI [9] CT(Lung) 1,018 scans (1,010 subjects) http://ncia.nci.nih.gov/ LUNA-16 888 scans https://luna16.grand-challenge.org/Data/ LUNA-DG [224] 887 scans https://github.com/meisun1207/LUNA-DG NLST [188, 189] 25,681 patients (77,040 images) https://cdas.cancer.gov/nlst/ COVID19-Diag 226 CT volumes https://github.com/MLMIP/COVID19-Diag ChestX-ray14 (NIH) [207] X-ray (Chest) 112,120 images (30,805 subjects) https://nihcc.app.box.com/v/ChestXray-NIHCC CheXpert (CXP) [85] 224,316 images (65,240 subjects) https://stanfordmlgroup.github.io/competitions/chexpert/ MIMIC-CXR (MMC) [90] 377,110 images (65,179 subjects) https://physionet.org/content/mimic-cxr/2.0.0/ PadChest [28] 160,000 images(67,000 subjects) https://bimcv.cipf.es/bimcv-projects/padchest/ Open-i dataset [46] 8,121 images http://openi.nlm.nih.gov/ Tawsifur [150, 38] 931 images https://www.kaggle.com/datasets/tawsifurrahman/covid19-radiography-database Skytells 1,017 images https://github.com/skytells-research/COVID-19-XRay-Dataset
Table 1: Public medical datasets used for generalization research
Dataset Modality (Organs) Number of cases Reference UK Biobank retinal photography dataset [185] Fundus photography (Eye) 58,700 patients https://www.ukbiobank.ac.uk/ EyePACs 88,702 images https://www.kaggle.com/c/diabetic-retinopathy-detection/data APTOS 3,662 images https://www.kaggle.com/c/aptos2019-blindness-detection Messidor 1,200 images https://www.adcis.net/en/third-party/messidor/ PALM 1,200 images https://palm.grand-challenge.org/ AV-DRIVE [82] 40 Images https://drive.grand-challenge.org/ LES-AV 22 images https://figshare.com/articles/dataset/LES-AV_dataset/11857698 HRF 45 images https://www5.cs.fau.de/research/data/fundus-images/ REFUGE 1,200 images https://refuge.grand-challenge.org/ REFUGE2 2,000 images https://refuge.grand-challenge.org/ STARE 20 images https://cecas.clemson.edu/~ahoover/stare/probing/index.html RIGA 750 images https://academictorrents.com/details/eb9dd9216a1c9a622250ad70a400204e7531196d DDR 13,673 images https://drive.google.com/drive/folders/1z6tSFmxW_aNayUqVxx6h6bY4kwGzUTEC RIMONEv2 [17] 455 images https://medimrg.webs.ull.es/ FGADR 1,842 images https://csyizhou.github.io/FGADR/ Endovis Challenge dataset [4] Endoscopic (Abdominal organs) 8 sequences https://endovissub2017-roboticinstrumentsegmentation.grand-challenge.org/ Heidelberg colorectal dataset [128] Laparoscopy (Colorectum) 30 laparoscopic videos https://robustmis2019.grand-challenge.org/Data/ CholecSeg8k [79] Laparoscopy (Abdomen) 17 videos from Cholec80 dataset http://camma.u-strasbg.fr/datasets SurgicalActions [164] Laparoscopy (Gynecologic organs) 160 videos http://www-itec.aau.at/ftp/datasets/SurgicalActions160 Cataract-101 [165] Video (Eye) 101 cataract surgeries http://ftp.itec.aau.at/datasets/ovid/cat-101/ PathVQA [76] Multiple modalities (multi-organs) 4,998 pathology images (multi-organs) https://github.com/UCSD-AI4H/PathVQA VQA-RAD[106] Radiology (multi-organs) 315 radiology images. https://osf.io/89kps/ OrganCMNIST CT (Abdomen) 23,583 https://medmnist.com/
Table 2: Public medical datasets used for generalization research
Domain generalization benchmarks
Zhang et al. [228] benchmarked [^2] https://github.com/MLforHealth/ClinicalDG the performance of eight DG methods on multi-site clinical times series from Intensive Care Units (ICUs) and chest X-ray imaging data from four sites. They compared the performances of these models to a baseline approach which consists of using Empirical Risk Minimization (ERM), where a single model is learned on pooled data across all training sources. In contrast to DG methods, this method assumes that the samples are IID. It minimizes the global average risk, defined as $R_{ERM}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(x_{i},y_{i};\theta)$ where $\mathcal{L}$ is the classification loss, $\theta$ is the model’s parameters, and $n$ is the number of samples [23].
In line with prior work on general imaging datasets [69], their experiments on real-world medical imaging data revealed that the current DG methods do not consistently achieve significant gains in OOD performance over ERM. In addition, the authors proposed a framework for augmenting clinical datasets via synthetic domain shifts and sampling bias.
Taxonomy of generalization research in medical classification domain
@@ 4.0.0.5 PANDA dataset
Depending on the assumed domain shift (covariate shift or concept shift), a plethora of methods have been proposed. To make it easier for the reader to find the methods suited to their problem, we have therefore chosen to first separate the methods based on this criterion. In this section, we present our categorization of methods based on covariate shift (Figure 1) and concept shift (Figure 2). These methods are detailed in the following sections (Section 6.1 and Section 6.2). Table 3 presents the notations used in this paper.
@@
$E,C$ Feature extractor (encoder)/classifier $H$ History of predictions
$F$ Projection layer $z^{c}$ Label prediction
@@
$\mathcal{S}$ Source domain $v$,$\delta$, $\beta$ Hyperparameters
@@
$M$ Number of source domains $\hat{z}^{f}$ Ensemble target of the features
$p$, $q$ Distribution after applying EMA to feature
$n_{i}$ Data size of source domain $i$ representations within previous epochs.
$n$ Data size of total training data $\hat{z}^{c}$ Ensemble target of the labels
R Risk function after applying EMA to label predictions
$y^{d}$ Label distribution within previous epochs.
$\hat{x}$ Augmented instance $N$ Total number of negative samples
@@
$\tau$ Temperature $\varrho$ Prototype
$\gamma$ Hyperparameter $I$ Individual regularization
@@
Figure 1: Literature survey tree for covariate shift
@@
Figure 2: Literature survey tree for concept shift and label noise
6.1 Covariate shift in medical imaging
Data heterogeneity is a key challenge for DL model generalizability. Covariate shift, in particular, is considered one of the most prominent shift in medicine. It is difficult to avoid this type of shift in medical imaging. It is mainly caused by the use of different type of acquisition systems and protocols, which may present notable differences among domains (i.e., changes in intensity values and contrast). Another factor to covariate shift is the differences in the characteristics of the lesions or diseases (shape, size, malignancy and location) and biological variations between patients (age, sex). Solutions for tackling the covariate shift can be categorized into: data manipulation (Section 6.1.1), representation (Section 6.1.2) and learning methods (Section 6.2.2).

### v3

@@ Abstract
Numerous Deep Learning (DL) classification models have been developed for a large spectrum of medical image analysis applications, which promises to reshape various facets of medical practice. Despite early advances in DL model validation and implementation, which encourage healthcare institutions to adopt them, a fundamental questions remain: how can these models effectively handle domain shift? This question is crucial to limit DL models performance degradation. Medical data are dynamic and prone to domain shift, due to multiple factors. Two main shift types can occur over time: 1) covariate shift mainly arising due to updates to medical equipment and 2) concept shift caused by inter-grader variability. To mitigate the problem of domain shift, existing surveys mainly focus on domain adaptation techniques, with an emphasis on covariate shift. More generally, no work has reviewed the state-of-the-art solutions while focusing on the shift types. This paper aims to explore existing domain generalization methods for DL-based classification models through a systematic review of literature. It proposes a taxonomy based on the shift type they aim to solve. Papers were searched and gathered on Scopus till 10 April 2023, and after the eligibility screening and quality evaluation, 77 articles were identified. Exclusion criteria included: lack of methodological novelty (e.g., reviews, benchmarks), experiments conducted on a single mono-center dataset, or articles not written in English. The results of this paper show that learning based methods are emerging, for both shift types. Finally, we discuss future challenges, including the need for improved evaluation protocols and benchmarks, and envisioned future developments to achieve robust, generalized models for medical image classification.
@@
Domain generalization , Medical image analysis , Covariate shift , Concept shift , Domain shift , Noisy labels
journal: arXiv
@@ Introduction
Deep Learning (DL) models are the current state-of-the-art method for medical image classification. The availability of high quality labeled data, typically through multi-site collaboration projects, has paved the way to employ these data driven-based approaches in supervised medical image analysis. Nowadays, DL models have achieved human level performances in different medical domains such as dermatology [1], oncology [2], histopathology [3] and ophthalmology [4].
@@
However, UDA methods are limited in practice, as they still require access to a part of the test-domain data during the training procedure. To overcome this limitation, Domain Generalization (DG) methods have emerged as a more promising solution. In DG, the goal is to develop a DL model that is able to generalize to one unseen target domain via learning from a single or multiple source domains, without having access to the testing data from the target domain. However, training DG methods using multi-source data (multi-DG) has been considered as costly since collecting medical data from multiple sources is challenging, and medical data are subject to privacy regulations. To address this problem, recent work focused on an additional research line, called single domain generalization (single-DG), in which the goal is to develop a DL model that is able to generalize to multiple target domains via learning from a single source domain [19]. Alternatively, semi-supervised domain generalization [20] combines the single-DG and multi-DG by using one labeled sources domain and several unlabeled source domain to boost the performances.
@@
DG in computer vision dataset is becoming an emerging field: numerous surveys have been proposed [21, 22]. In the medical field, research has focused on domain adaptation [23] or unsupervised domain adaptation [24]. Other medical research has reviewed the problem of learning with noisy labels [25, 26, 27]. However, to the best of our knowledge, no medical review has studied the problem of generalization of DL models in the medical field with a focus on both domain shift problems: covariate shift and concept shift. A study of the current DL methods tackling these problems is thus necessary for guiding practitioners and researchers in understanding the challenges and existing trends in the field. In particular, exploring and analyzing these methods would help identify the limits and the best methods. This would lead to more efficient and robust DL systems, enabling a broader applicability of AI in different environment healthcare settings. This paper presents the first systematic review of generalization research in medical image classification. It aims to answer the following Research Questions (RQ):
  • RQ1: What are the state-of-the-art methods in medical image classification targeting domain shift in the literature?
    Significance: A taxonomy and a clustering of similar methods would help analyze the performances for different shift types. It would also help identifying which generalization techniques are most effective under different circumstances.
  • RQ2: What are the related areas in which generalization research can be applied?
    Significance: Identifying related areas will help understand the scenarios where this research can be combined with other studies or applied in practice.
  • RQ3: What are the best practices for implementing generalization techniques in research?
    Significance: Identifying open-source libraries and implementations details would enhance generalization research in the medical domain.
  • RQ4: What are the key challenges and future promises for generalization research?
    Significance: Identifying key challenges and future research areas with potential for significant advancements in generalization research is crucial for guiding researchers toward the most promising directions.
In this paper, we make the following key contributions:
  • We present the first systematic survey on generalization research for medical image analysis based on covariate shift and concept shift. Based on the assumed shift type, the reader can refer to methods in our taxonomy.
  • We present public medical datasets and open-source libraries to enhance future research in this field.
  • We study the recent trends in generalization research and found that learning-based methods are showing an increased interest. In particular, foundation models hold promises for enhanced generalizability.
  • Our analysis shows that this research is applied to wide areas in medical imaging, including: X-ray, fundus photography, dermoscopic imaging, and pathology. There is a need for benchmarking strategies to better assess these methods.
@@
The organization of this paper is as follows. In Section 3, we briefly describe the problem of domain generalization. In Section 4, we introduce our methodology for literature review. In Section 5, we present our taxonomy, in which we review DL methods that have dealt with covariate shift in the medical domain (Section 5.1) and DL methods aiming to overcome the problem of concept shift and noisy labels (Section 5.2). In Section 6, we present public medical datasets used for generalization research. Section 7 discusses the benefit of current DG methods based on the results of challenge data. Furthermore, it presents trends in DG development, related research to DG, implementation details, and future directions. Finally, Section 8 concludes this work.
@@
Research Methodology
To address our research questions posed, the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guiding principles for conducting systematic reviews [28] was applied to select papers which develop solutions in generalization research. Figure 1 displays the PRISMA flowchart conducted in this work. The review process consisted of gathering studies using Scopus database. The search strategy was piloted by one reviewer using the following query: “domain generalization”  OR “noisy labels” OR “covariate shift” OR “concept shift”. This search was done within Article title, Abstract, and keywords. We included papers published from 1 January 2020 to 10 April 2023 (included). A total of 2086 papers were found. First, the search results were reviewed and duplicated records were removed using Zotero. This resulted in 2027 papers. Abstracts and titles were manually reviewed. Papers were included if they were dealing with medical image classification and deep learning methods. They were excluded if they met one of the following criteria: 1) the paper was not accessible in English, 2) the paper was a review, 3) the paper was a result of a challenge, 4) the paper was a benchmark, 5) the paper included exactly one dataset provided it is not a multi-center dataset, 6) the paper did not propose a new method for tackling concept shift or covariate shift. When in doubt about the eligibility of the study, the full text was retrieved ant reviewed. The total number of papers considered in this survey was 77 papers.
We developed a data extraction form comprising different items related to the research questions. It included the following items: 1) title of the article, 2) year of publishing, 3) modality, 4) organ, 5) task, 6) dataset, 7) type of shift: covariate or concept shift, 8) deep model: the deep learning technique used in the study, 9) code availability and 10) dataset availability. One reviewer collected data from each report.
For a fair comparison, we have chosen to report the results of papers using the same testing subset.
Figure 1: PRISMA flowchart for systematic review of generalizing methods.
What are the state-of-the-art methods in medical image classification targeting domain shift in the literature?
@@
The identified papers were reviewed and a taxonomy was proposed based on the common methodology and the assumed shift they solve. Depending on the assumed domain shift (covariate shift or concept shift), a plethora of methods have been proposed. To make it easier for the reader to find the methods suited to their problem, we have therefore chosen to first separate the methods based on this criterion (Figure 2). In this section, we present our categorization of methods based on covariate shift (Figure 3) and concept shift (Figure 4). These methods are detailed in the following sections (Section 5.1 and Section 5.2). Table 1 presents the notations used in this paper.
@@
$E,C$ Feature extractor (encoder)/classifier $I$ Individual regularization
$M$ Number of source domains $z^{c}$ Label prediction
@@
$\mathcal{S}$ Source domain $v$, $\delta$, $\beta$, $\gamma$ Hyperparameters
@@
$P$ Total number of positive samples $N$ Total number of negative samples
$p$, $q$ Distribution R Risk function
$n_{i}$ Data size of source domain $i$ $n$ Data size of total training data
$\hat{x}$ Augmented instance $y^{d}$ Label distribution
@@
$\tau$ Temperature
@@
Figure 2: The generalization taxonomy proposed in our domain generalization research for medical image classification. The motivation of this work is based on four research questions. Depending on the domain shift type, 3 methods of categories were identified for covariate shift and concept shift.
@@
classification (Section 5.1), basic, l sep=10mm, [Data manipulation (Section 5.1.1), root, l sep=10mm, [Data homogenization [29] [30] [31] [32] [33] (Section 5.1.1), xnode, l sep=10mm, ] [Data augmentation [19] [20] [32] [34] [35] [36] [37] [38] [39] [40] [41] [42] [43] [44] (Section 5.1.1) , xnode, l sep=10mm, ] ] [Representation
learning (Section 5.1.2), root, l sep=10mm, [Adversarial learning [38] [45] [46] [47] [48] (Section 5.1.2), xnode, l sep=10mm, ] [Feature distribution alignment [49] (Section 5), xnode, l sep=10mm, ] [Contrastive learning [50] [51] (Section 5), xnode, l sep=10mm, ] [ Consistency regularization [20] [52] [53] [54] [55] (Section 5), xnode, l sep=10mm, ] ] [Learning
strategies (Section 5.1.3), root, l sep=10mm, [Ensemble learning [35] [56] [57] [58] (Section 5.1.3), xnode, l sep=10mm, ] [Test-time augmentation [41] [59] (Section 5.1.3), xnode, l sep=10mm, ] [Incremental [60] [61] (Section 5.1.3), xnode, l sep=10mm, ] [Self-supervised learning [39] [43] [61] [62] (Section 5.1.3), xnode, l sep=10mm, ] [Meta-learning [63] [64] [65] (Section 5.1.3), xnode, l sep=10mm,] [Gradient operation [66] (Section 5.1.3), xnode, l sep=10mm, ] [Distributionally Robust optimization [59] (Section 5.1.3), xnode, l sep=10mm, ] [Multi-task learning [67] [68] [69] (Section 5.1.3), xnode, l sep=10mm, ] ] ]
Figure 3: Literature survey tree for covariate shift.
@@
(Section 5.2.2), root, l sep=10mm, [Distance metric learning [85] [86] [87] [88] (Section 5.2.2), xnode, l sep=10mm, ] [ Active learning [70](Section 5.2.2), , xnode, l sep=10mm, ] [Zero-shot learning [89] (Section 5.2.2), xnode, l sep=10mm, ] [ Gradient [90] (Section 5.2.2), , xnode, l sep=10mm,] [Robust loss function [76] [78] [80] [91] [92] [93] [94] [95] [96] (Section 5.2.2), xnode, l sep=10mm, ] [Graph [84] [93] [95] [97] (Section 5.2.2), xnode, l sep=10mm, ] [Meta-model [98] (Section 5.2.2), xnode, l sep=10mm, ] ] [Collaborative
methods (Section 5.2.3), root, l sep=10mm, [Co-training [99] [100] (Section 5.2.3), xnode, l sep=10mm, ] [Co-teaching [76] [80] [81] [101] [102](Section 5.2.3), xnode, l sep=10mm, ] [Knowledge distillation [103] (Section 5.2.3), xnode, l sep=10mm, ]] ]
Figure 4: Literature survey tree for concept shift and label noise.
5.1 Covariate shift in medical image classification
Data heterogeneity is a key challenge for DL model generalizability. Covariate shift, in particular, is considered one of the most prominent shift in medicine. It is difficult to avoid this type of shift in medical imaging. It is mainly caused by the use of different type of acquisition systems and protocols, which may present notable differences among domains (i.e., changes in intensity values and contrast). Another factor to covariate shift is the differences in the characteristics of the lesions or diseases (shape, size, malignancy and location) and biological variations between patients (age, sex). Solutions for tackling the covariate shift can be categorized into: data manipulation (Section 5.1.1), representation learning (Section 5.1.2) and learning methods (Section 5.2.2).